OpenStudy (anonymous):

Solve for x in the proportion . x = 8 and x = −8 x = 8 and x = −15 x = −8 x = −3 and x = −5

OpenStudy (uri):

Question?

OpenStudy (uri):

wut.

OpenStudy (uri):

wot.

thomaster (thomaster):

You need to solve for x but you didn't post the problem

OpenStudy (austinl):

There has to be more to the question, or you are trolling us.....

OpenStudy (uri):

Well there is no question just answer choices..

OpenStudy (anonymous):

the question is 2x/16x+30=-1/x

OpenStudy (austinl):

\[\frac{2x}{16x+30}=-\frac{1}{x}\] Is this correct?

OpenStudy (anonymous):

yes

OpenStudy (austinl):

I would cross multiply first.

OpenStudy (anonymous):

alright .

OpenStudy (anonymous):

i got 2x^2=-16x-30. that is right so far right

OpenStudy (austinl):

Yep.

OpenStudy (austinl):

Then move the right side over to the left side.

OpenStudy (anonymous):

so it would be 2x^2+16x+30 right?

OpenStudy (austinl):

Right! Then you would factor that.

OpenStudy (anonymous):

so i would factor out a GCF so that it becomes smaller?

OpenStudy (austinl):

Indeed!

OpenStudy (anonymous):

the GCF is 2 so then i would get x^2+8x+15=x

OpenStudy (austinl):

Right, so what would the factor be?

OpenStudy (anonymous):

2?

OpenStudy (austinl):

No, I mean, factor it out. a(x+b)(x+c) a=2 right? so what are b and c?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

@austinL this should help http://static.someecards.com/someecards/usercards/1340050882014_3068657.png

OpenStudy (austinl):

No. What times what is 15, but when added is 8?

OpenStudy (anonymous):

3 and 5

OpenStudy (austinl):

Correct. There you go... However, I believe we made an error somewhere. It should be -3 and -5.

OpenStudy (anonymous):

ya i was about to say that because there are no answers that say 3 and 5

OpenStudy (anonymous):

thank you austinL!!! ^_^

OpenStudy (austinl):

No problem.